Thu Jan 15 19:29:31 PST 2004 — Brief web service outage. We experienced a problem with our web server cluster this evening. During the brief (5 minute) time of the outage web services were not available. Operations restored service and are investigating the cause. -Sonic.net Operations
IPv6 Tunnel Outage.
Wed Jan 14 21:15:56 PST 2004 — IPv6 Tunnel Outage. A misconfiguration on the system which provides customer IPv6 tunnels caused it to lose connectivity to the IPv4 Internet. The configuration has been repaired and the tunneling system is operating normally. -Nathan
Graton Rooftop Outage.
Wed Jan 14 09:14:18 PST 2004 — Graton Rooftop Outage. A switch serving Graton Rooftop customers has failed. We are working with an on-site tech to resolve the issue. -Matt
Stockton area DSL services degraded.
Tue Jan 13 20:01:48 PST 2004 — Stockton area DSL services degraded. Our upstream Internet transit provider for our Stockton facility is having routing issues. There is currently no estimated time to repair. DSL customers in that area are experiencing increased latency (high ping times). -John, Nathan and Kelsey
Update Tue Jan 13 20:37:34 PST 2004 — Our provider reports the problem fixed, and we confirm that. -Sonic NOC
Focal router problems: The router at Focal in
Mon Jan 12 06:15:50 PST 2004 — Focal router problems: The router at Focal in San Francisco is currently down. All dialup numbers ending in 9606, 9811 and 1001 are returning authentication failures. An employee is currently en-route to investigate the issue and should arrive in approximately 45 minutes.
As always, alternate dialup numbers are available. To locate a dialup in your area, visit:
-Nathan and John
Mon Jan 12 08:11:32 PST 2004 — Update: Service is restored.
Graton Rooftop Outage.
Fri Jan 9 08:21:00 PST 2004 — Graton Rooftop Outage. A router serving Graton Rooftop customers has failed. A replacement is being prepared. ETR is about 30 minutes. -Matt
Fri Jan 9 08:32:50 PST 2004 — Correction – This issue is in Graton, not Occidental.
Web server issue.
Wed Jan 7 08:53:44 PST 2004 — Web server issue. One of our web servers stopped serving requests for a few minutes this morning. This caused some intermittent unavailability of some web pages. The issue has been resolved. -Matt and Geoff
Router maintenance Wednesday morning.
Tue Jan 6 17:27:36 PST 2004 — Router maintenance Wednesday morning. We are doing a board swap at the direction of Cisco in the router at Focal between 5 and 6 AM Wednesday 1/7/04 to correct ongoing problems. The outage should be 15 minutes or less and should only be noticeable as a couple of brief interruptions on some routes. -John, Zeke and Nathan
Update — Besides the interruptions mentioned, there will also be an outage for customers using the Focal dialup numbers while the router is out of service. Other dialup numbers will be unaffected. -John and Eli
Update: Wed Jan 7 05:19:40 PST 2004 — Preparations are taking longer than expected. The actual outage time will now be closer to 6:30. -John and Zeke
Linux Kernel Upgrades: We are in the process…
Mon Jan 5 12:46:04 PST 2004 — Linux Kernel Upgrades: We are in the process of upgrade all of our Linux kernels to the latest stable release in order to fix a potential local exploit. All customer accessible servers except for our shell server have already been upgraded. Bolt.sonic.net will be upgraded in a few minutes. -Kelsey
Media server change done.
Mon Jan 5 17:57:01 PST 2004 — Media server change done. We completed the migration to the new media server for Streaming RealMedia. This was a transparent migration, so if you notice any issues with streaming content of your website, please let our support department know. -Kevan